If F varies jointly as q^2 and f=-24 when q=2 and h=-2, find K
Answer by venugopalramana(3286)

If F varies jointly as q^2 and f=-24 when q=2 and h=-2, find K
PLEASE TYPE QUESTION CORRECTLY..I HOPE YOUR QUESTION IS
If F varies jointly as q^2......AND.....H.... and f=-24 when q=2 and h=-2, find K
ASSUMING SO
F=KHQ^2
-24=K*-2*2^2=-8K
8K=24
K=3